what is the poh of a 7.2 x 10-8 m blood sample

1
Expert's answer
2022-05-20T03:40:02-0400

[H+] = 7,2 * 10-8

1. first we find the pH: pH = -lg [ H+ ]

pH = -lg7,2 *10-8 =8 - lg7,2 =8 - 0,857 =7,143

2. the sum of pH and poh in any solution equal is 14. Using this, we find pOH:

pH + pOH = 14

pOH = 14 - pH = 14 - 7,143 = 6,857

ANSWER: pOH = 6.857